OpenStudy (anonymous):

Calculus - Related Rates: From a certain spot X, a car travels due north at a speed of 70 miles per hour. Ten miles east of this spot X, another car travels due south at a speed of 80 miles per hour. How fast is the distance between the two cars increasing the moment they are 40 miles apart?

OpenStudy (zarkon):

differentiate \[x^2+y^2=z^2\] with respect to time t

OpenStudy (anonymous):

2x dx/dt + 2y dy/dt =2z dz/dt

OpenStudy (zarkon):

good...did you draw a picture...what is x ,y, and z

OpenStudy (zarkon):



OpenStudy (zarkon):

we can then make a large triangle...

OpenStudy (zarkon):



OpenStudy (zarkon):

I guess I shouldn't have used X since they are referring to it as a point

OpenStudy (anonymous):

i'm with you so far

OpenStudy (zarkon):

y=10,z=40...find x

OpenStudy (anonymous):

x= \[\sqrt{1500}\]

OpenStudy (anonymous):

sorry dude, this script is killing my ie8. had to bail and switch to mozilla.

OpenStudy (zarkon):

do you know what x' and y' are?

OpenStudy (anonymous):

yes, x' is 70mph and y' is 80mph

OpenStudy (zarkon):

no

OpenStudy (zarkon):

y is a fixed 10 miles...so y'=0

OpenStudy (anonymous):

ah hah

OpenStudy (zarkon):

and x' is not 70

OpenStudy (zarkon):

north and south are both vertical displacements so x'=70+80=150

OpenStudy (anonymous):

ok, gotcha

OpenStudy (zarkon):

plug in all your numbers into your derivative and solve for z'

OpenStudy (anonymous):

145.24mph?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

awesome, thank you for your help. I was doing some crazy stuff I didn't need to do.

OpenStudy (zarkon):

np